OpenStudy (anonymous):

for rational function f, identify any vertical asymptotes. f(x) = (4-x)/(x+5)

OpenStudy (anonymous):

@Easyaspi314 Could you help me

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

-5 is one

OpenStudy (anonymous):

x = -5 is a V.A. (thats where denominator is zero)

OpenStudy (anonymous):

y = -1 is H.A.

OpenStudy (anonymous):

Becuase the limit of the function as x------>infinity is -1.

OpenStudy (anonymous):

Thank you! I have another one. For the rational function f(x) = -100(2x-1)(4x+3)(x-2) / x(x-1)(x-2). Identify the horizontal asymptote.

OpenStudy (anonymous):

I think its y = -800?

OpenStudy (anonymous):

Yes, you are correct!

OpenStudy (anonymous):

For rational functoin f(x) = x^2 + x + 6 / x^2 - 10x + 24, identify any vertical asymptotes.

OpenStudy (anonymous):

Denominator: (x-4)(x-6), hence x = 4 and x= 6 Note: numerator is not factorable, so no cancellation there.

OpenStudy (anonymous):

So the vertical asymptotes are x = 4 and x = 6?

OpenStudy (anonymous):

yes